This should be super easy, but I am not doing something right, however I cannot figure out what it is. I know that the answer is 100 (from creating a table), but I need to show my work algebraically. When I do it algebraically I can't get the right answer for X. Please help!!!!
$8(X)=$200+$10(X-40)

Answer by josgarithmetic(39617)

EDIT: THIS solution contained a mistake, now been fixed!
8X=200+10(X-40), dropping the dollar symbol, focusing just on the numbers without money unit;
8X=200+10X-400, Distributive Property
0=200-400+2X, Additive Inverse of 8X and Commutative Property for Addition
2X-200=0, (.......what's my reason?)
2X=200,
X=100, X=100